A pure titanium cube has an edge length of 2.68 . How many titanium atoms does it contain? Titanium has a density of 4.50 g/cm3
Answer #1
volume of Ti = 2.68^3 = 19.2 cm^3
mass of Ti = volume x density = 19.2 x 4.50 = 86.6 g
atomic mass of Ti = 47.9
moles of Ti = mass / atomic mass = 86.6/47.9 = 1.81 mol
1 moles of Ti contains 6.02 x 10^23 Ti atoms
Thus,
number of Ti in the cube = 1.81 x 6.02 x 10^23 = 1.09 x 10^24
Answer #2
2.68 in x 2.54 = 6.8072 cm

Volume cube = ( 6.8072)^3 =351.4 cm^3

mass = d x V = 4.50 g/cm^3 x 351.4 cm^3 = 1419.44 g

moles Ti = 1419.44 g/ 47.867 g/cm^3= 29.65

atoms = 29.65 x 6.02 x 10^23 =1.78 x 10^25